### v1

# Comprehensive Analysis of Relative Pressure Estimation Methods Utilizing 4D Flow MRI

Abstract
4D flow MRI allows for the estimation of three-dimensional relative pressure fields, providing rich pressure information, unlike catheterization and Doppler echocardiography, which provide one-dimensional pressure drops only. The accuracy of one-dimensional pressure drops derived from 4D flow has been explored in previous literature, but additional work must be done to evaluate the accuracy of three-dimensional relative pressure fields. This work presents an analysis of three state-of-the-art relative pressure estimators: virtual Work-Energy Relative Pressure (vWERP), the Pressure Poisson Estimator (PPE), and the Stokes Estimator (STE). Spatiotemporal behavior and sensitivity to noise were determined in silico. Estimators were validated with a type B aortic dissection (TBAD) flow phantom with varying tear geometry and an array of twelve catheter pressure measurements. Finally, the performance of each estimator was evaluated across eight patient cases. In silico pressure field errors were lower in STE compared to PPE, although PPE pressures were less affected by noise. High velocity gradients and low spatial resolution contributed most significantly to local variations in 3D error fields. Low temporal resolution leads to highly transient peak pressure events being averaged, systematically underestimating peak pressures. In the flow phantom analysis, vWERP was the most accurate method, followed by STE and PPE. Each pressure estimator strongly correlated with ground truth pressure values despite the tendency to underestimate peak pressures. Patient case results demonstrated that the pressure estimators could be feasibly integrated into a clinical workflow.
{IEEEkeywords}
Relative pressure estimation, hemodynamics, 4D flow MRI, aortic dissection
1 Introduction
[…]
Relative pressure differences are important when assessing the severity of a variety of cardiovascular diseases, including aortic stenosis [1], aortic coarctation [2], mitral valve regurgitation [3], and aortic dissection [4]. Until recently, relative pressure differences could only be evaluated by invasive catheterization, which carries inherent risks as a diagnostic tool [5], or by Doppler echocardiography [6]. While Doppler echocardiography is a useful diagnostic tool in a subset of disorders, it measures only one velocity direction. It is, therefore, limited to simple pressure models, such as the simplified Bernoulli equation. 4D flow MRI [7] allows for the measurement of three-dimensional velocity fields over the cardiac cycle, capturing rich blood flow information. This rich flow information enables the estimation of additional hemodynamic variables, such as wall shear stress, kinetic energy, and relative pressure [8, 9, 10]. Pressure estimators that utilize 4D flow MRI have the potential to increase the diagnostic information available by mapping relative pressure across various cardiovascular domains. One-dimensional pressure drops derived from catheter measurements or Doppler echocardiography have been in use for decades, but 4D flow MRI holds the potential for utilizing three-dimensional pressure fields as a diagnostic tool.
We will herein focus on analyzing the performance of pressure estimators in the context of type B aortic dissection (TBAD) [11]. In brief, TBAD refers to a partial thickness entry tear in the descending aorta, creating a second pathway for blood known as the false lumen (FL). The false lumen often reconnects with the true lumen (TL) distally through an exit tear. TBAD contains a wide variety of complex flow patterns and challenging geometries, making it a valuable setting for the evaluation of pressure estimators. TBAD contains regions of low velocities, high velocities, narrow geometries, and highly transient pressure events. Low-velocity regions are helpful when analyzing sensitivity to noise. Narrow domains are essential when analyzing spatial resolution limits; highly transient events are valuable when analyzing temporal resolution limits. Finally, high spatial gradients have been reported to be particularly challenging when estimating hemodynamic parameters [12, 13], meriting further study.
Pressure estimators are highly dependent on the image quality of a given 4D flow acquisition [14]. Consequently, it is crucial to stress-test these estimators under a variety of imaging conditions. Previous work has been done to validate and explore the use of various pressure estimators [15, 16, 17, 18, 14, 19], but additional analysis is necessary. In particular, the spatial distribution of errors requires further study. This work compares CFD-derived pressure fields with estimated pressure fields in a novel voxel-to-voxel analysis. The spatial errors resulting from velocity field corruptions are also determined. Flow phantoms with 12-point arrays of pressure catheter measurements are utilized as ground truth in an in vitro study, enhancing our understanding of the spatial variation of errors in pressure estimations derived from 4D flow acquisitions. In sum, these analyses will add to our understanding of the spatial variations in pressure estimator accuracy rather than the accuracy of one-dimensional pressure drops.
In this paper, the relative pressure estimators Virtual Work-Energy Pressure (vWERP), Pressure Poisson Estimator (PPE), and Stokes Estimator (STE) are analyzed across in silico, in vitro, and in vivo TBAD datasets. For the in silico analysis, we utilize patient-specific CFD simulations to generate synthetic 4D flow MRI as input. The resulting estimated relative pressure fields are compared directly to CFD-derived relative pressure fields. Spatiotemporal and noise sensitivity analyses are performed for both the pressure fields and pressure drops. For the in vitro analysis, we utilize 4D flow acquired across three TBAD flow phantom variations and an array of ground truth pressure catheter measurements [20]. Variations in the flow phantoms allow for a qualitative analysis of the changing pressure fields as the TBAD false lumen tears are modified. Validation of the pressure estimators is performed via comparison to catheter pressure measurements. Finally, pressure drops estimated in eight patient cases are compared across methods for agreement. Sources of error are identified, and each pressure estimator is compared to determine their pros and cons in different use cases.
2 Material and Methods
2.1 Relative Pressure Estimation Algorithms
Relative pressure estimation algorithms utilize full-field 4D flow MRI velocity measurements as input. These algorithms can be split into fully spatial and plane-to-plane estimators. Fully spatial estimators produce a four-dimensional (3D + time) relative pressure field and include PPE and STE. Plane-to-plane estimators produce a time-resolved relative pressure drop between two user-defined planes (e.g., the pressure drop from the aortic root to the FL exit tear over one cardiac cycle) and include vWERP. The pressure estimation algorithms used herein are derived from the incompressible Navier-Stokes equations [21] given by (1) and (2)
$\rho\frac{\partial{\bf\it u}}{\partial t}+\rho({\bf\it u}\cdot\nabla){\bf\it u}$ $=-\nabla p+\mu\Delta{\bf\it u}\quad\text{on }\Omega$ (1)
$\nabla\cdot{\bf\it u}$ $=0\quad\text{on }\Omega,$ (2)
where $\Omega$ denotes the spatial domain. Equation (1) can be rearranged to estimate pressure gradients by plugging in the measured velocity field, ${\bf\it u}_{m}$.
$R_{m}=-\rho\left(\frac{\partial{\bf\it u}_{m}}{\partial t}+({\bf\it u}_{m}\cdot\nabla){\bf\it u}_{m}\right)+\mu\nabla^{2}{\bf\it u}_{m}\quad\text{on }\Omega$ (3)
2.1.1 Pressure Poisson Estimator (PPE)
The pressure Poisson estimator [14, 22] is derived by taking the divergence of (3), simplifying using the assumption of incompressible flow (2), and subsequently solving the resulting Poisson equation with Neumann boundary conditions.
$\nabla^{2}\hat{p}$ $=-\rho\nabla\cdot({\bf\it u}_{m}\cdot\nabla{\bf\it u}_{m})\quad\text{on }\Omega$ (4)
$\nabla\hat{p}\cdot{\bf\it n}$ $=-\rho({\bf\it u}_{m}\cdot\nabla{\bf\it u}_{m})\cdot{\bf\it n}\quad\text{on }\partial\Omega,$ (5)
where $\partial\Omega$ is the boundary of the spatial domain and $\nabla\hat{p}$ is the estimated pressure gradient field.
2.1.2 Stokes Estimator (STE)
The Stokes estimator [23, 24] utilizes a divergence-free virtual field, $\tilde{{\bf\it u}}$, as a regularizer (6).
$\nabla\hat{p}-\mu\nabla^{2}\tilde{{\bf\it u}}$ $=R_{m}\quad\text{on }\Omega$ (6)
$\nabla\cdot\tilde{{\bf\it u}}$ $=0\quad\text{on }\Omega$ (7)
$\tilde{{\bf\it u}}$ $=0\quad\text{on }\partial\Omega$ (8)
It does not require higher-order derivatives and has been reported to handle boundaries better because it does not prescribe boundary conditions directly on the pressure field [15], as PPE does.
2.1.3 Virtual Work-Energy Relative Pressure (vWERP)
Virtual work-energy relative pressure is derived from virtual work-energy equations and is based on the principle of conservation of energy. The virtual work-energy equation is given by multiplying (1) by a virtual field, ${\bf\it w}$, and integrating over the domain $\Omega$ between two planes.
[…]
The virtual velocity field ${\bf\it w}$ is constructed by solving a Stokes flow problem over the user-specified flow domain. After integrating and rearranging (9), (the details of which can be found in [19]), the pressure drop between two planes is given by
$\Delta p=\frac{-1}{Q}\left(\frac{\partial}{\partial t}K_{e}+A_{e}+V_{e}\right)$ (10)
where $K_{e}$ denotes the kinetic energy held within the blood flow, $A_{e}$ denotes the rate of energy entering/leaving the system, and $V_{e}$ is the rate of energy dissipation due to viscous forces. $Q$ is the volumetric flow rate between the two planes over which the pressure drop is calculated. vWERP has shown substantial improvement over older plane-to-plane estimators, such as the unsteady Bernoulli equation [25] and the original WERP [19].
2.2 Implementation of Relative Pressure Estimation Methods
STE, PPE, and vWERP were all implemented using in-house MATLAB codes. All methods utilized a midpoint temporal discretization scheme and second-order spatial derivatives where applicable. The relevant equations for STE and vWERP were discretized with a finite difference method (FDM) that used the measured velocity values directly as input [19]. PPE was discretized using a second-order accurate finite volume method (FVM). For PPE, an FVM was chosen over an FDM to reduce errors introduced by boundary normal vector estimation when prescribing Neumann boundary conditions (5) on the pressure field [26]. Finally, the velocity input was nearest-neighbor upsampled depending on the resolution. The 3.0, 2.0, and 1.5 mm images were upsampled by factors of 4, 4, and 2, respectively. This led to an effective resolution of 0.75, 0.50, and 0.75 mm. All velocity gradients, laplacians, and temporal derivatives were calculated using the raw velocity input before upsampling.
2.3 In Silico Analysis
2.3.1 Model Generation
A patient-specific CFD model of a TBAD patient was utilized as a test bed to analyze PPE, STE, and vWERP. The CFD model allowed for the creation of synthetic 4D flow images at different spatiotemporal resolutions and noise levels to analyze the pressure estimation methods comprehensively. Model geometry was sourced from a clinical CT, segmented in 3D Slicer [27], and meshed with 1.1 mm elements and a 0.9 mm boundary layer in Simmetrix SimModeler [28]. Boundary conditions were prescribed using a 4D flow MRI scan (3T; voxel size = ${\sim}1.5\text{ mm}\times 1.5\text{ mm}\times 2.5\text{ mm}$; temporal resolution = 47 ms, VENC = 2.0 m/s) of the same patient. Inflow and outflow boundary conditions were imposed at the aortic root, brachiocephalic artery, left common carotid artery, left subclavian artery, and true lumen outlet based on flow rates interpolated from 4D flow to 1.0 ms resolution. Flow resistance was estimated by measuring the area of each outlet. The aortic wall was assumed to be completely rigid. A single cardiac cycle was simulated in $\mathcal{C}$Heart [29]. The solved pressure field, $p_{\text{CFD}}$, was projected from the unstructured CFD mesh onto isotropic 0.50 mm and 0.75 mm grids at a 20 ms sampling rate. These projections served as the ground truth.
Figure 1: Overview of in silico model at peak systolic flow. (a) Solved velocity field on unstructured FEM mesh. (b) Solved velocity field after projection to a voxelized domain. Planes used for pressure drop analysis are shown in blue and gold. (c) Noisy velocity magnitude images at tested SNR levels. Non-fluid regions are set to zero to aid comparison.
2.3.2 Spatiotemporal Analysis
The solved velocity field, ${\bf\it u}_{\text{CFD}}$, was projected from the unstructured CFD mesh onto a variety of regular, isotropic grids to mimic MRI image format (Fig. 1ab). Projections were made onto 1.5, 2.0, and 3.0 mm grids at 20 ms, 40 ms, and 60 ms sampling rates for a total of nine spatiotemporal combinations. This allowed for a spatiotemporal analysis of PPE, STE, and vWERP to evaluate the robustness of each pressure estimation method across a variety of resolutions. Performance was analyzed by propagating each of the ${\bf\it u}_{\text{CFD}}$ fields through each estimator and directly comparing the estimated pressure results ($p_{\text{PPE}}$, $p_{\text{STE}}$) against the ground truth pressure solution, $p_{\text{CFD}}$. Since vWERP is a plane-to-plane estimator, vWERP-derived pressure drops were compared directly with pressure drops sourced from $p_{\text{CFD}}$. Nine unique planes (for eight separate pressure drops) were generated for this purpose (Fig. 1b).
As fully spatial relative pressure estimation methods, PPE and STE could be compared directly with $p_{\text{CFD}}$. Relative error was evaluated at each timestep with (11) and across the entire time domain with (12).
$\epsilon_{\text{rel}}(t)=\frac{\left\lVert p_{\text{CFD}}(t)-\hat{p}(t)\right\rVert_{0,2}}{\left\|p_{\text{CFD}}\right\|_{0,\infty}}$ (11)
$\varepsilon_{\text{rel}}=\frac{\left\lVert p_{\text{CFD}}-\hat{p}\right\rVert_{0,2}}{\left\lVert p_{\text{CFD}}\right\rVert_{0,2}}$ (12)
RMSE was also calculated over the entire spatial domain at each timestep and across the entire spatiotemporal domain. In addition to these error metrics, linear regression plots were also generated.
vWERP pressure drop estimates were compared directly with pressure drops selected from $p_{\text{CFD}}$. Planes were drawn in ParaView [30]. Pressure drops selected from $p_{\text{STE}}$ and $p_{\text{PPE}}$ were compared with pressure drops selected from $p_{\text{CFD}}$ as well. As in the fully spatial evaluation, linear regression plots were generated.
2.3.3 Noise Sensitivity Analysis
To assess the impact of estimate accuracy, noise fields [31] were added to the projected velocity fields to evaluate the performance of PPE, STE, and vWERP. The noise addition methodology was as follows: Velocity fields were first converted to phase data by specifying a synthetic VENC that was 1.1 times the maximum velocity component across the entire time-resolved velocity field. Synthetic magnitude data was created by multiplying the voxelized CFD mask by a non-zero constant (leaving the non-fluid region equal to 0). The synthetic magnitude and phase images were combined to create complex images that were subsequently transformed to k-space via the fast Fourier Transform. Zero-mean, complex Gaussian noise was added to k-space. The variance of the complex Gaussian was defined to achieve a specified SNR using the equation $\text{SNR}_{\text{db}}=10\log{\frac{P_{x}}{P_{n}}}$ ), where $P_{x}$ is the power of the signal and $P_{n}$ is the power of the noise (which is equivalent to the variance for white Gaussian noise). Finally, the k-space data was transformed back to image space via the inverse Fourier Transform, and the phase images were converted back to velocity images using the previously specified VENC. Twenty-five realizations of SNR = 10 noise fields and SNR = 30 noise fields (Fig. 1c) were calculated for a total of 50 realizations per resolution, resulting in 450 realizations in total. The SNR = $\inf$ was performed implicitly during the spatiotemporal analysis. The same error metrics used in the previous spatiotemporal analysis were used for the noise sensitivity analysis, with some modifications to the correlation analysis — instead of performing a linear regression over one pressure field, linear regressions were performed over all 25 noise realizations at once. Finally, error averages, minimums, and maximums were calculated for each spatiotemporal and noise combination (e.g., the average error was calculated over the 25 realizations of the 1.5 mm $\times$ 40 ms and SNR = 30 combination). Finally, pressure drops were evaluated for their robustness to noise with (13) and (14).
$e_{\Sigma}=\frac{\left\lVert\Delta p_{\text{original}}(t)-\Delta p_{\text{corrupted}}(t)\right\rVert_{0,1}}{\left\lVert\Delta p_{\text{original}}(t)\right\rVert_{0,1}}$ (13)
$e_{\text{max}}=\frac{\left\lVert\Delta p_{\text{original}}(t_{n})-\Delta p_{\text{corrupted}}(t_{n})\right\rVert_{0,\infty}}{\left\lVert\Delta p_{\text{original}}(t_{n})\right\rVert_{0,\infty}}$ (14)
2.4 Flow Phantoms
PPE, STE, and vWERP were validated with an in vitro flow setup [20]. 4D flow MRI was acquired, and twelve pressure catheter measurements per TBAD tear variation served as ground truth.
2.4.1 Flow Setup
A full description of the flow phantom can be found in [20] (pump model, exact flow specifications, etc.). A brief overview is provided here. The flow phantom geometry was designed as an embedded compliant 3D-printed TBAD derived from a clinical CT exam. Physiologically accurate pulsatile flow and pressure boundary conditions were applied. 4D flow MRI (3T; voxel size = $1.5\times 1.5\times 1.5$ mm; temporal resolution = 50 ms; VENC = 1.2–1.7 m/s) was acquired across three phantom variations: one baseline phantom (TBADOR) and two additional variations with identical anatomy but with either a reduced size FL entry tear (TBADENT) or a reduced size FL exit tear (TBADEXT) Fig. 2bc. Twelve catheter pressure measurements were acquired at consistent locations across each TBAD tear variation, for a total of 36 unique measurements (Fig. 2a).
Figure 2: Overview of in vitro flow set up. (a) Catheter measurement points and cut planes across TBAD tear variations. (b) Entry tears across all TBAD tear variations. (c) Exit tears across all TBAD tear variations.
2.4.2 Validation Methodology
Eleven time-resolved pressure drops derived from the twelve pressure catheter locations serve as the ground truth. To validate PPE and STE, the entire pressure field was calculated ($p_{\text{STE}},p_{\text{PPE}}$), and individual pressure drops were found by spatially averaging the pressure values at each catheterization plane (i.e., the specified inlet and outlet planes) and calculating the difference at each time point to produce temporally-resolved, one-dimensional pressure drops. vWERP, being a plane-to-plane estimator, natively calculated these one-dimensional pressure drops, so no additional post-processing was required. Linear regressions were performed to analyze the performance of each pressure estimator.
2.4.3 Corruption Analysis
As with the CFD model, the velocity fields were corrupted prior to pressure estimation in order to evaluate each method’s robustness to noise. Noise, however, is already present in the acquired 4D flow, so any additional perturbations are referred to as “corruptions”. Each velocity component ($u_{x},u_{y},u_{z}$) was corrupted with a zero-mean, truncated Gaussian [32] with standard deviation dependent on the level of additional corruption desired ($\Delta\text{SNR}$) according to equation (15).
$\sigma_{v}=\frac{\sqrt{2}u_{\max}}{\pi\cdot\Delta\text{SNR}}$ (15)
Each TBAD tear variation was evaluated over 50 realizations with $\Delta$SNR = 30 and $\Delta$SNR = 10, for a total of 300 corrupted realizations per pressure estimation method. Pressure drop deviations were calculated for each unique corrupted velocity field realization. This included measures of mean deviation (13) and max deviation (14). Error averages, minimums, and maximums were recorded across each of the 50 realization combos (e.g., $\text{TBAD}_{\text{ENT}}$ at $\Delta$SNR = 10) and are reported in the supplement.
2.5 Patient Cases
The final analysis utilized eight patient cases (Fig. 3ab) to assess the clinical feasibility of pressure estimation methods and to qualitatively evaluate pressure trends against what is known in the literature, including our knowledge of FL growth rate [33]. 4D flow was acquired on 3T MRI scanners (MR750, GE Medical Systems, Milwaukee, WI, USA; Ingenia, Philips, Best, Netherlands) with the following scan parameters: voxel size = ${\sim}1.5\text{ mm}\times 1.5\text{ mm}\times 2.5\text{ mm}$, average temporal resolution = 47 ms, VENC = 2.0 m/s. Agreement between the pressure estimators was also evaluated. 4D flow was acquired from a previous IRB-approved study [34] (HUM00120679). The data was segmented by thresholding patient MRA data and subsequently interpolating the resulting segmentation onto the 4D flow image space (Fig. 3cd). Inlet and outlet planes were generated in a 4D flow MATLAB toolbox [35]. FL GR was measured clinically as the change in maximum aortic diameter. Each pressure estimator was run on all eight in vivo cases. The pressure fields calculated by the fully spatial pressure estimators were analyzed for their coherence. The 4D pressure fields were also converted to pressure drops (inlet–TL outlet and inlet–FL outlet) to allow for comparisons with vWERP.
Figure 3: Two representative patient cases. (a) High-growth patient case MRA with 3D segmentation overlaid. The TL segmentation is blue, and the FL segmentation is gold. (b) Low-growth patient case MRA with 3D segmentation overlaid. (c) High-growth patient case 4D flow magnitude with velocity streamlines at peak systole overlaid. (d) Low-growth patient case 4D flow magnitude with velocity streamlines at peak systole overlaid.
3 Results
3.1 Evaluation of In Silico TBAD Model
3.1.1 Spatiotemporal Analysis
Although the spatiotemporal analysis was performed on nine unique spatiotemporal resolutions, only the 1.5 mm $\times$ 40 ms results are visualized in Fig. 4. Additional correlation plots and pressure drop visualizations can be found in section 1 of the supplement. Fig. 4abc presents the CFD, PPE, and STE pressure fields, respectively, at peak systole. Relative error values with no added noise are recorded in Table 1. Fig. 4f shows the correlation between the pressure fields calculated by STE, PPE, versus the CFD ground truth. The estimated pressure fields were highly consistent — both $p_{\text{STE}}$ and $p_{\text{PPE}}$ were very strongly correlated with $p_{\text{CFD}}$. PPE consistently mildly underestimated the ground truth pressure compared to STE.
Fig. 4e (top) shows the pressure drop from the aortic inlet to the TL outlet as calculated by vWERP, STE, and PPE. Table 2 summarizes the pressure drop linear regression results for the spatiotemporal analysis. All correlation coefficients were near unity for the 1.5 mm and 2.0 mm images but degraded dramatically for the 3.0 mm images. PPE correlation coefficients decreased the most dramatically for the 3.0 mm cases. Additional press drop linear regression plots can be found in subsection 1.5 of the supplement.
Figure 4: Overview of the in silico 1.5 mm $\times$ 40 ms results. (a) CFD ground truth pressure field at peak systole. (b) STE pressure field at peak systole. Aortic inlet (i) and true lumen outlet (TL4) are marked with arrow for subfigure d. (c) PPE pressure field at peak systole (d) Pressure drop from aortic inlet to TL outlet. (e) Linear regression of same pressure drop. (f) Linear regressions performed on STE and PPE pressure fields across all timesteps. The heatmaps indicate the distribution of pressure values in each 2D bin. The number of points in each 2D bin is normalized to the number of points in the entire column. The 1D histogram above the heatmap correlation plots illustrates the number of points present within each column.
Table 1: In Silico Total Relative Error ($\varepsilon_{\text{rel}}$) for Fully Spatial Pressure Fields
STE PPE
20 ms 40 ms 60 ms 20 ms 40 ms 60 ms
1.5 mm 3.33% 8.76% 16.89% 8.30% 11.98% 16.69%
2.0 mm 19.57% 18.34% 24.83% 6.79% 8.51% 13.69%
3.0 mm 78.98% 80.43% 85.87% 101.76% 103.10% 110.15%
Table 2: In Silico Pressure Drop Slopes and Correlations
vWERP STE PPE
20 ms 40 ms 60 ms 20 ms 40 ms 60 ms 20 ms 40 ms 60 ms
1.5 mm $m$ 0.86 0.85 0.92 0.99 0.92 1.04 0.94 0.93 0.99
$r^{2}$ 1.00 0.99 0.98 1.00 0.99 0.98 1.00 0.99 0.97
2.0 mm $m$ 0.89 0.88 0.96 1.09 1.08 1.17 1.01 0.99 1.06
$r^{2}$ 0.98 0.98 0.99 0.98 0.98 0.99 0.99 0.99 0.99
3.0 mm $m$ 0.36 0.35 0.36 0.54 0.52 0.54 0.42 0.41 0.40
$r^{2}$ 0.36 0.34 0.30 0.39 0.37 0.33 0.21 0.19 0.16
3.1.2 Sensitivity Analysis Results
The sensitivity analysis for the 1.5 mm $\times$ 40 ms resolution is summarized in Fig. 5. The relative error over the cardiac cycle (11) is shown in Fig. 5a for STE and PPE. Although both methods had similar maximum errors, STE was slightly more accurate on average. Correlation heatmaps for the fully spatial analysis and the plane-to-plane pressure drop analysis at SNR = 10 are shown in Fig. 5b and d, respectively. A pressure error map for an SNR = 10 1.5 mm $\times$ 40 ms case is shown in Fig. 5c. Errors were lowest in the distal true lumen. In general, errors grew when moving proximally from the true lumen outlet. The exception to this trend was in the extremely narrow region of the proximal true lumen, which is shown in more detail in Fig. 5c (right). The range of total relative errors (12) is shown in Table 3. Averaged total relative errors can be found in subsection 1.3 of the supplement. Relative errors for STE were lower than for PPE; however, STE generally had larger ranges of errors, indicating that it was more sensitive to noise.
Table 3: In Silico Total Relative Error ($\varepsilon_{\text{rel}}$) Range for Sensitivity Analysis
STE PPE
SNR dx/dt 20 ms 40 ms 60 ms 20 ms 40 ms 60 ms
30 1.5 mm 2.08% 1.74% 1.48% 1.61% 2.08% 2.34%
2.0 mm 3.43% 4.48% 5.20% 2.72% 2.52% 3.14%
3.0 mm 8.04% 12.50% 16.18% 11.75% 12.97% 16.90%
10 1.5 mm 1.28% 3.24% 3.37% 2.76% 4.00% 4.20%
2.0 mm 7.06% 5.14% 5.69% 2.59% 2.78% 6.15%
3.0 mm 23.97% 28.18% 30.06% 15.42% 16.99% 28.98%
In general, all methods were resilient to noise; only the 3.0 mm pressure traces in supplement subsection 1.6 had a noticeable spread of pressure values due to corrupted velocity fields. Furthermore, the calculated deviation values indicated that the amount of deviation from the baseline (no noise) cases increased substantially for the 3.0 mm cases. Across all resolutions and methods, the deviation values for the 1.5 mm and 2.0 mm cases remained below 4% in almost all cases and frequently below 2%. Of the 3.0 mm cases, the vWERP estimations showed the highest deviations from the baseline case, followed by STE, and then PPE a distant third place.
Figure 5: Overview of in silico 1.5 mm $\times$ 40 ms sensitivity analysis. (a) Relative error over cardiac cycle. (b) Correlation heatmaps for the fully spatial analysis over all 25 noise realizations. (c) Pressure error map (d) Correlation heatmaps for the plane-to-plane pressure drop analysis across all 25 noise realizations. The heatmaps indicate the distribution of pressure values in each 2D bin. The number of points in each 2D bin is normalized to the number of points in the entire column. The 1D histogram above the heatmap correlation plots illustrates the number of points present within each column.
3.2 Flow Phantom Analysis
Figure 6: Overview of in vitro flow phantom analysis. (a) TBADENT velocity field at peak systole. (b) TBADENT STE pressure field at peak systole. (c) TBADENT PPE pressure field at peak systole. (d) Linear regression across all models and planes with no corruption. (e) Example pressure trace over the cardiac cycle. The pressure drop was calculated from the model inlet to the model outlet (i to o in Fig. 2). The “x” markers indicate where 4D flow timeframes were present. (f) Correlation heatmaps for each method. Linear regressions were run over all 50 $\Delta$SNR=10 corruption realizations per TBAD tear variation for a total of 150 unique datasets per correlation heatmap. The heatmaps indicate the distribution of pressure values in each 2D bin. The number of points in each 2D bin is normalized to the number of points in the entire column. The 1D histogram above the heatmap correlation plots illustrates the number of points present within each column.
3.2.1 Validation of Pressure Estimators
A broad overview of the flow phantom analysis results is given in Fig. 6, with TBADENT chosen as the representative TBAD tear variation. More extensive results, including those for TBADOR and TBADENT, can be found in the supplement. Qualitatively, the pressure fields estimated by STE and PPE demonstrated the features one would expect from observing the TBADENT velocity field (Fig. 6a). For instance, the high pressure impingement zone immediately following the entry tear in Fig. 6bc closely resembles the impingement zone in the corresponding FSI model of TBADENT, shown in Fig. 4c of Zimmermann et al. [20]. Both the STE (Fig. 6b) and PPE (Fig. 6c) pressure fields exhibited a large pressure drop across the FL entry tear, an increase in pressure where the FL entry tear jet impinged on the wall of the FL, and higher pressure drops from the aortic inlet (i) to each TL measurement point (TL 1–5) compared to each FL measurement point (FL 1–5). The TBADOR and TBADEXT pressure fields were similarly coherent with our expectations. Fig. 6df includes data from all TBAD tear variations. Across all TBAD tear variations, vWERP, STE, and PPE were strongly correlated ($r^{2}=0.91$) with catheter measurements (Fig. 6d), with a tendency to underestimate peak pressures (Fig. 6e). This trend holds across all pressure traces, which can be explored in further detail in subsection 2.3 of the supplement.
3.2.2 Corruption Analysis
Fig. 6f illustrates the effect of the $\Delta$SNR = 10 corruption across all 50 noise realizations for each pressure estimation method. Corruption of the velocity fields mildly degraded the average accuracy, while the correlations were moderately weaker ($r^{2}=0.69$). The pressure drop mean deviation (13) and max deviation (14) results for TBADENT are summarized in Table 4. For the FL and TL rows, deviations were averaged across all numbered pressure drops for that lumen (e.g., aortic inlet to TL1, aortic inlet to TL2, etc.). Additional tables for TBADOR and TBADEXT can be found in the supplement in subsection 2.2, and a more in-depth analysis for each pressure trace and the range of estimations can be found in the supplement subsection 2.3. As expected, the average deviation increased with noise. Average deviations were also higher in the true lumen across all models. PPE was significantly more resilient to corruption compared to vWERP and STE — the average PPE deviations were roughly half of the other method’s deviations.
Table 4: In Vitro Average Deviations Across TBADENT
vWERP STE PPE
SNR = 30 SNR = 10 SNR = 30 SNR = 10 SNR = 30 SNR = 10
$e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$
FL 1.52% 1.45% 5.41% 4.73% 1.20% 1.22% 4.42% 3.83% 0.57% 0.60% 1.81% 1.85%
TL 2.95% 2.27% 12.69% 8.90% 2.76% 2.15% 10.80% 7.72% 1.38% 1.11% 4.49% 3.50%
Outlet 2.38% 2.50% 8.55% 8.25% 1.22% 1.16% 4.73% 4.15% 0.58% 0.57% 1.88% 1.75%
3.3 Patient Cases
For all eight patient cases, PPE and STE produced pressure fields that were coherent with our expectations — relative pressure was highest in the aortic arch and steadily decreased when moving distally. Pressure drops were visibly higher in the TL, where most of the net flow was present [36]. Pressure drops in the FL were generally lower than in the TL, although the magnitude of this trend was dependent on the calculated growth rate of each patient [33]. As expected, lower FL growth rate patients had a higher pressure drop across the FL, and higher FL growth rate patients had a lower pressure drop across the FL (Fig. 7ab). The PPE, STE, and vWERP pressure drops were analyzed for agreement in Fig. 7cd, and additional comparisons can be found in section 3 of the supplement. Correlations between all pressure estimation methods were extremely strong ($r^{2}$ = 0.97–0.98). Pressure drops calculated by vWERP tended to be the highest, followed by STE ($m$ = 0.95–0.97 when compared to vWERP), and finally, PPE ($m$ = 0.87–0.93 compared to vWERP). This ordering of the results matched the outcome of the in vitro phantom analysis, where vWERP also produced the highest pressure drops.
Standard practice was to run the patient cases with nearest-neighbor upsampling by a factor of 2. At this level, pressure estimations were completed in a matter of minutes as timed by MATLAB’s timeit function. Timing results for a representative patient case were: 527.4 s for STE, 151.4 s for PPE, 192.7 s for vWERP when calculating the FL pressure drop, and 41.3 s for vWERP when calculating the TL pressure drop.
Figure 7: In vivo analysis results. (a) STE pressure field of a high-growth patient case at peak systole. (b) STE pressure field of a low-growth patient case at peak systole. (c) TL and FL correlation plot of STE vs. vWERP across all patient cases. (d) TL and FL correlation plot of PPE vs. vWERP across all patient cases.
4 Discussion
The pressure estimators vWERP, STE, and PPE were evaluated across in silico, in vitro, and in vivo TBAD datasets. The in silico analysis was performed on a CFD model with varying spatiotemporal resolution and synthetic noise levels. Notably, the in silico analysis allowed for comparisons between estimated pressure fields and the ground truth CFD solution. This allowed for a novel analysis of the spatial distributions of errors. The in vitro analysis was performed on a compliant TBAD flow phantom with three unique entry and exit tear variants. Twelve pressure catheter measurements were made in each flow phantom tear variation, combining the benefits of acquired 4D flow and significantly more ground truth pressure measurements than would be allowed clinically. The in vivo analysis was performed on eight patient cases. While no ground truth was present, this analysis allowed for a qualitative view of pressure fields and a chance to observe the relative pressure estimators in a clinical context.
4.1 Spatial Error Distribution
An important novelty of the fully spatial in silico analysis was the ability to precisely analyze the spatial variation of errors in the estimated pressure fields. In general, errors were highest on the vessel wall boundaries and in areas with high velocity gradients. These two factors combined to become particularly impactful in the 3.0 mm in silico images — an extremely narrow region in the proximal TL had high velocities, leading to sharp velocity gradients and extensive partial voluming. Thus, absolute errors of over 15 mmHg occurred and relative errors of over 100% occurred, as seen in Fig. 8b and Table 1. To further corroborate that partial voluming/interpolation was causing high errors, an additional analysis was performed by translating the CFD model by 1/3 of a given voxel size. Each translation forced different interpolations of the boundary, changing TL pressure drops and relative error. One can imagine a 6 mm diameter region that is either more accurately represented by two 3 mm voxels or poorly represented after a 1 mm translation, which causes the same data to be represented over three voxels with partial voluming.
In the 1.5 mm and 2.0 mm cases, the TL pressure drop was resilient to shifting the model; however, the 3.0 mm case gave drastically different pressure drops depending on the shift, ranging from 10.9 mmHg to 18.2 mmHg as the maximum pressure drop. Because the relative pressures were calculated in reference to the TL outlet, errors in the proximal TL shifted the rest of the estimated pressure field by a constant for both STE and PPE. This behavior leads to high errors when compared to the ground truth CFD data (and when compared to the TL pressure drop ground truth), but not when compared to the FL pressure drop ground truth. Because the rest of the model proximal to the narrow TL was disrupted by the same constant, the four FL pressure drops we probed remained accurate.
In the patient cases, small segments of 2–8 voxels with erroneous velocity measurements led to small regions (2–8 voxels as well) with massive pressure errors. High (but localized) erroneous velocity gradients led to corresponding localized pressure errors. In response to these findings, care was taken to pre-process patient data to remove 4D flow voxels with magnitudes/orientations significantly different from their neighbors. If STE and PPE are used clinically, care must be taken to appropriately filter out erroneous velocity measurements to ensure errors are not magnified in the corresponding relative pressure fields. Errors generally increased when moving proximally from the TL outlet, which is expected given that the relative pressure fields were shifted in reference to the TL outlet pressure.
Figure 8: Overview of in silico origin shift analysis for STE. (a) Overview of origin shift method for one voxel in 2D. Each square indicates a different shift in the origin by a subvoxel amount, forcing a different interpolation. (b) Example of the large errors in the FL and proximal TL after and unfavorable shit. (c) Relative errors over the cardiac cycle for five different origin shifts. Temporal resolution was fixed to 60 ms for each variation in spatial resolution.
4.2 Effect of Resolution
Increased spatial resolution generally increased the accuracy of the estimated pressure fields. Accuracy dipped when temporal resolution decreased mainly because the pressure estimators did not have the time frames available to capture the highly transient pressure drop at peak systole. Errors were primarily caused by underestimating pressure extrema (both systolic peaks and diastolic valleys) and a temporal shift in the latter half of the cardiac cycle. This behavior was more drastic in the 4D flow MRI acquisitions compared to the synthetic 4D flow images. The synthetic 4D flow images were constructed with instantaneous velocity measurements, which meant there was no temporal averaging. Temporal averaging, of course, occurred in the 4D flow scans. In the in silico analysis, the estimators may have missed peak pressures, but they accurately estimated the pressures based on the time frames they had access to. Meanwhile, the 4D flow scans both undersampled the peak systolic pressure drop and temporally averaged the time frames that were sampled, leading to more pronounced underestimation.
4.3 Effect of Noise
In agreement with Bertoglio et al. [14], the average errors across all noisy realizations were roughly equivalent to the cases with no added noise, experimentally confirming that vWERP, STE, and PPE are unbiased estimators. Whereas higher resolution was almost always better in the SNR=inf cases, this trend did not precisely hold in the noisy cases. Low spatial resolution combined with high temporal resolution resulted in a higher spread in estimated pressures, although the average pressures across noise realizations were still close to the SNR=inf cases. Deviations from the baseline pressure drop estimation were also routinely higher in the TL across both in silico and in vitro analyses.
4.4 Pros and Cons of Pressure Estimators
Each pressure estimator recovered flow phantom ground truth pressures with a minimum correlation coefficient of 0.9 and a tendency to mildly underestimate peak pressure drops, albeit with varying severity. More clear differences in accuracy were present, however, in the sensitivity and spatiotemporal analyses. Based on our analysis, vWERP should be the preferred estimator when calculating one-dimensional pressure drops. It had the highest accuracy in the 4D flow MRI acquisitions and was significantly faster than STE. Although vWERP performed worse than PPE and STE in the in silico cases, vWERP’s improved performance in the in vitro cases suggests that there are aspects of MRI simulation that we are missing in our synthetic MRI generation methodology. When a fully 3D pressure field is desired, the ideal estimator depends on the SNR and resolution of the 4D flow acquisition. STE should be the standard choice (especially at lower resolutions). However, if a low SNR is expected, PPE may be better than STE. Similar to [15], STE outperformed PPE in terms of accuracy; however, PPE appeared to have a slight edge in its resistance to noise. PPE’s additional resistance to noise may be due to the slightly different implementation of PPE, which was implemented as an FVM rather than an FDM. It should be noted that Nolte et al. [15] used a different implementation of PPE and STE, utilizing backward differences for temporal derivatives (rather than midpoint) and an FEM. Although PPE was faster than STE (roughly 2.5 minutes vs. 8.7 minutes of CPU time), both methods calculated pressure fields in a reasonable amount of time.
4.5 Limitations and Future Work
A synthetic MRI generation protocol accepted in the past was utilized [31] but could be improved [37, 38]. Future work would involve integrating more advanced synthetic MRI modeling to determine results in hypothetical clinical scenarios more accurately. The CFD model design and synthetic image generation procedure were both reasonably straightforward. This decision was made because the projection to a regular grid required that an extremely large number of elements (roughly 2.6 million) be present to preserve higher-order derivatives when projecting to the finer grid resolutions. Although this model was sufficiently representative, future work could involve an FSI model with compliant walls. Finally, the in vivo analysis did not include patient catheter data, although such invasive data is not typically acquired (particularly in the FL) in patients with aortic dissection.
5 Conclusion
In this work, we analyzed vWERP, STE, and PPE over in silico, in vitro, and in vivo datasets. Of the fully spatial estimators, STE was shown to be more accurate overall, whereas PPE was shown to be more resilient to noise. vWERP did not perform as well as either STE or PPE in the in silico analysis. In the in vitro analysis, STE continued to outperform PPE in accuracy but was mildly less resilient to corruption of the velocity field. vWERP moderately outperformed both STE and PPE in accuracy. PPE remained the most resilient to noise. Finally, the pressure estimators were evaluated qualitatively and for agreement over patient cases. PPE and STE agreed strongly, and vWERP tended to estimate the largest pressure drop, mimicking the trend of the in vitro analysis. All pressure estimators were remarkably resilient to noise and successfully estimated the ground truth pressure, albeit with tendencies to underestimate. The time to compute the pressure estimations was on the order of minutes, demonstrating the possibility of being integrated into a clinical workflow.

### v2

@@
-Comprehensive Analysis of Relative Pressure Estimation Methods Utilizing 4D Flow MRI
+Comprehensive Analysis of Relative Pressure Estimation Methods Utilizing 4D-Flow MRI
 Abstract
-4D flow MRI allows for the estimation of three-dimensional relative pressure fields, providing rich pressure information, unlike catheterization and Doppler echocardiography, which provide one-dimensional pressure drops only.
-The accuracy of one-dimensional pressure drops derived from 4D flow has been explored in previous literature, but additional work must be done to evaluate the accuracy of three-dimensional relative pressure fields.
+Magnetic resonance imaging (MRI) can estimate three-dimensional (3D) time-resolved relative pressure fields using 4D-flow MRI, thereby providing rich pressure field information.
+Clinical alternatives include catheterization and Doppler echocardiography, which only provide one-dimensional pressure drops.
+The accuracy of one-dimensional pressure drops derived from 4D-flow has been explored previously, but additional work is needed to evaluate the accuracy of 3D relative pressure field estimates.
 This work presents an analysis of three state-of-the-art relative pressure estimators: virtual Work-Energy Relative Pressure (vWERP), the Pressure Poisson Estimator (PPE), and the Stokes Estimator (STE).
-Spatiotemporal behavior and sensitivity to noise were determined in silico.
-Estimators were validated with a type B aortic dissection (TBAD) flow phantom with varying tear geometry and an array of twelve catheter pressure measurements.
+The spatiotemporal characteristics and sensitivity to noise were determined in silico.
+Estimators were then validated using a type B aortic dissection (TBAD) flow phantom with varying tear geometry and twelve catheter pressure measurements.
 Finally, the performance of each estimator was evaluated across eight patient cases.
-In silico pressure field errors were lower in STE compared to PPE, although PPE pressures were less affected by noise.
-High velocity gradients and low spatial resolution contributed most significantly to local variations in 3D error fields.
-Low temporal resolution leads to highly transient peak pressure events being averaged, systematically underestimating peak pressures.
+In silico pressure field errors were lower in STE compared to PPE, although PPE pressures were less noise sensitive.
+High velocity gradients and low spatial resolution contributed most significantly to local variations in 3D pressure field errors.
+Low temporal resolution lead to systematic underestimation of highly transient peak pressure events.
 In the flow phantom analysis, vWERP was the most accurate method, followed by STE and PPE.
-Each pressure estimator strongly correlated with ground truth pressure values despite the tendency to underestimate peak pressures.
-Patient case results demonstrated that the pressure estimators could be feasibly integrated into a clinical workflow.
+Each pressure estimator was strongly correlated with ground truth pressure values, despite the tendency to underestimate peak pressures.
+Patient case results demonstrated that each pressure estimator could be feasibly integrated into a clinical workflow.
 {IEEEkeywords}
-Relative pressure estimation, hemodynamics, 4D flow MRI, aortic dissection
+Relative pressure estimation, pressure gradient, hemodynamics, 4D-flow MRI, aortic dissection
 1 Introduction
@@
 Relative pressure differences are important when assessing the severity of a variety of cardiovascular diseases, including aortic stenosis [1], aortic coarctation [2], mitral valve regurgitation [3], and aortic dissection [4].
-Until recently, relative pressure differences could only be evaluated by invasive catheterization, which carries inherent risks as a diagnostic tool [5], or by Doppler echocardiography [6].
-While Doppler echocardiography is a useful diagnostic tool in a subset of disorders, it measures only one velocity direction.
-It is, therefore, limited to simple pressure models, such as the simplified Bernoulli equation.
-4D flow MRI [7] allows for the measurement of three-dimensional velocity fields over the cardiac cycle, capturing rich blood flow information.
-This rich flow information enables the estimation of additional hemodynamic variables, such as wall shear stress, kinetic energy, and relative pressure [8, 9, 10].
-Pressure estimators that utilize 4D flow MRI have the potential to increase the diagnostic information available by mapping relative pressure across various cardiovascular domains.
-One-dimensional pressure drops derived from catheter measurements or Doppler echocardiography have been in use for decades, but 4D flow MRI holds the potential for utilizing three-dimensional pressure fields as a diagnostic tool.
-We will herein focus on analyzing the performance of pressure estimators in the context of type B aortic dissection (TBAD) [11].
-In brief, TBAD refers to a partial thickness entry tear in the descending aorta, creating a second pathway for blood known as the false lumen (FL).
+Until recently, relative pressure differences could only be evaluated by invasive catheterization, which carries inherent risks as a minor surgery [5], or noninvasively by Doppler echocardiography [6].
+While Doppler echocardiography is a useful diagnostic tool in a subset of disorders, it can fail in some patients and only measures one velocity direction.
+Therefore, it is limited to simple pressure models, such as the simplified Bernoulli equation.
+4D-flow MRI [7] measures three-dimensional (3D), time-resolved velocity fields throughout the cardiac cycle, capturing rich blood flow information.
+This flow information enables the estimation of additional hemodynamic variables, such as wall shear stress, kinetic energy, and relative pressure [8, 9, 10].
+Pressure estimators that utilize 4D-flow MRI have the potential to increase the diagnostic information available by mapping relative pressure across various cardiovascular domains.
+One-dimensional pressure drops derived from catheter measurements or Doppler echocardiography have been in use for decades, but 4D-flow MRI holds the potential for utilizing four-dimensional pressure (3D + time) fields as a diagnostic tool.
+Herein, we focus on analyzing the performance of pressure estimators in type B aortic dissection (TBAD) [11].
+TBAD refers to a partial thickness entry tear in the descending aorta, creating a second pathway for blood within a false lumen (FL).
 The false lumen often reconnects with the true lumen (TL) distally through an exit tear.
 TBAD contains a wide variety of complex flow patterns and challenging geometries, making it a valuable setting for the evaluation of pressure estimators.
-TBAD contains regions of low velocities, high velocities, narrow geometries, and highly transient pressure events.
+TBAD contains regions with narrow geometries, highly transient pressure events, and low velocities immediately adjacent to high velocities.
+Narrow domains are essential when analyzing spatial resolution limits; highly transient events are valuable when analyzing temporal resolution limits.
 Low-velocity regions are helpful when analyzing sensitivity to noise.
-Narrow domains are essential when analyzing spatial resolution limits; highly transient events are valuable when analyzing temporal resolution limits.
-Finally, high spatial gradients have been reported to be particularly challenging when estimating hemodynamic parameters [12, 13], meriting further study.
-Pressure estimators are highly dependent on the image quality of a given 4D flow acquisition [14].
-Consequently, it is crucial to stress-test these estimators under a variety of imaging conditions.
-Previous work has been done to validate and explore the use of various pressure estimators [15, 16, 17, 18, 14, 19], but additional analysis is necessary.
-In particular, the spatial distribution of errors requires further study.
-This work compares CFD-derived pressure fields with estimated pressure fields in a novel voxel-to-voxel analysis.
-The spatial errors resulting from velocity field corruptions are also determined.
-Flow phantoms with 12-point arrays of pressure catheter measurements are utilized as ground truth in an in vitro study, enhancing our understanding of the spatial variation of errors in pressure estimations derived from 4D flow acquisitions.
+Finally, high spatial velocity gradients are particularly challenging when estimating hemodynamic metrics [12, 13], meriting further study.
+Pressure estimators are highly dependent on the image quality of a given 4D-flow acquisition [14].
+Consequently, it is crucial to stress-test these estimators by perturbing and modifying various image properties (e.g., spatial resolution) and determining their impact on pressure estimation accuracy.
+Previous work has validated and explored the use of various pressure estimators [15, 16, 17, 18, 14, 19], but additional analysis is necessary – the work above has focused on the accuracy of one-dimensional pressure drops, which don’t take full advantage of the rich velocity information provided by 4D-flow.
+Furthermore, the spatial distribution of errors requires further study.
+Herein, we compare CFD-derived pressure fields with estimated pressure fields in a novel voxel-to-voxel analysis.
+The spatial distribution of errors resulting from synthetic noise is also determined.
+Compliant TBAD flow phantoms with 12-point arrays of pressure catheter measurements are utilized as ground truth in an in vitro study, enhancing our understanding of the spatial variation of errors in pressure estimations derived from 4D-flow acquisitions.
 In sum, these analyses will add to our understanding of the spatial variations in pressure estimator accuracy rather than the accuracy of one-dimensional pressure drops.
-In this paper, the relative pressure estimators Virtual Work-Energy Pressure (vWERP), Pressure Poisson Estimator (PPE), and Stokes Estimator (STE) are analyzed across in silico, in vitro, and in vivo TBAD datasets.
-For the in silico analysis, we utilize patient-specific CFD simulations to generate synthetic 4D flow MRI as input.
+In this paper, we analyzed three relative pressure estimators: Virtual Work-Energy Pressure (vWERP), the Pressure Poisson Estimator (PPE), and the Stokes Estimator (STE), across in silico, in vitro, and in vivo TBAD datasets.
+For the in silico analysis, we utilize patient-specific CFD simulations to generate synthetic 4D-flow MRI as input.
 The resulting estimated relative pressure fields are compared directly to CFD-derived relative pressure fields.
 Spatiotemporal and noise sensitivity analyses are performed for both the pressure fields and pressure drops.
-For the in vitro analysis, we utilize 4D flow acquired across three TBAD flow phantom variations and an array of ground truth pressure catheter measurements [20].
+For the in vitro analysis, we utilize 4D-flow acquired across three different compliant TBAD flow phantom with varying FL entry and FL exit tears combined with an array of ground truth pressure catheter measurements [20].
 Variations in the flow phantoms allow for a qualitative analysis of the changing pressure fields as the TBAD false lumen tears are modified.
@@
 2.1 Relative Pressure Estimation Algorithms
-Relative pressure estimation algorithms utilize full-field 4D flow MRI velocity measurements as input.
+Relative pressure estimation algorithms utilize full-field 4D-flow MRI velocity measurements as input.
 These algorithms can be split into fully spatial and plane-to-plane estimators.
 Fully spatial estimators produce a four-dimensional (3D + time) relative pressure field and include PPE and STE.
-Plane-to-plane estimators produce a time-resolved relative pressure drop between two user-defined planes (e.g., the pressure drop from the aortic root to the FL exit tear over one cardiac cycle) and include vWERP.
+Plane-to-plane estimators produce a time-resolved relative pressure drop between two user-defined planes (e.g., the pressure drop from the aortic root to the FL exit tear throughout one cardiac cycle) and include vWERP.
 The pressure estimation algorithms used herein are derived from the incompressible Navier-Stokes equations [21] given by (1) and (2)
@@
 The pressure Poisson estimator [14, 22] is derived by taking the divergence of (3), simplifying using the assumption of incompressible flow (2), and subsequently solving the resulting Poisson equation with Neumann boundary conditions.
-$\nabla^{2}\hat{p}$ $=-\rho\nabla\cdot({\bf\it u}_{m}\cdot\nabla{\bf\it u}_{m})\quad\text{on }\Omega$ (4)
-$\nabla\hat{p}\cdot{\bf\it n}$ $=-\rho({\bf\it u}_{m}\cdot\nabla{\bf\it u}_{m})\cdot{\bf\it n}\quad\text{on }\partial\Omega,$ (5)
+$\nabla^{2}\hat{p}$ $=\nabla\cdot R_{m}\quad\text{on }\Omega$ (4)
+$\nabla\hat{p}\cdot{\bf\it n}$ $=R_{m}\cdot{\bf\it n}\quad\text{on }\partial\Omega,$ (5)
 where $\partial\Omega$ is the boundary of the spatial domain and $\nabla\hat{p}$ is the estimated pressure gradient field.
@@
 The Stokes estimator [23, 24] utilizes a divergence-free virtual field, $\tilde{{\bf\it u}}$, as a regularizer (6).
-$\nabla\hat{p}-\mu\nabla^{2}\tilde{{\bf\it u}}$ $=R_{m}\quad\text{on }\Omega$ (6)
+$\nabla\hat{p}-\nabla^{2}\tilde{{\bf\it u}}$ $=R_{m}\quad\text{on }\Omega$ (6)
 $\nabla\cdot\tilde{{\bf\it u}}$ $=0\quad\text{on }\Omega$ (7)
@@
 2.2 Implementation of Relative Pressure Estimation Methods
-STE, PPE, and vWERP were all implemented using in-house MATLAB codes.
+STE, PPE, and vWERP were all implemented using in-house MATLAB code.
 All methods utilized a midpoint temporal discretization scheme and second-order spatial derivatives where applicable.
@@
 2.3.1 Model Generation
-A patient-specific CFD model of a TBAD patient was utilized as a test bed to analyze PPE, STE, and vWERP.
-The CFD model allowed for the creation of synthetic 4D flow images at different spatiotemporal resolutions and noise levels to analyze the pressure estimation methods comprehensively.
+A patient-specific CFD model of a TBAD patient was used as a test bed to analyze PPE, STE, and vWERP.
+The CFD model allowed for the creation of synthetic 4D-flow images at different spatiotemporal resolutions and noise levels to analyze the pressure estimation methods comprehensively.
 Model geometry was sourced from a clinical CT, segmented in 3D Slicer [27], and meshed with 1.1 mm elements and a 0.9 mm boundary layer in Simmetrix SimModeler [28].
-Boundary conditions were prescribed using a 4D flow MRI scan (3T; voxel size = ${\sim}1.5\text{ mm}\times 1.5\text{ mm}\times 2.5\text{ mm}$; temporal resolution = 47 ms, VENC = 2.0 m/s) of the same patient.
-Inflow and outflow boundary conditions were imposed at the aortic root, brachiocephalic artery, left common carotid artery, left subclavian artery, and true lumen outlet based on flow rates interpolated from 4D flow to 1.0 ms resolution.
+Boundary conditions were prescribed using a 4D-flow MRI scan (3T; voxel size = ${\sim}1.5\text{ mm}\times 1.5\text{ mm}\times 2.5\text{ mm}$; temporal resolution = 47 ms, VENC = 2.0 m/s) of the same patient.
+Inflow and outflow boundary conditions were imposed at the aortic root, brachiocephalic artery, left common carotid artery, left subclavian artery, and true lumen outlet based on flow rates interpolated from 4D-flow to 1.0 ms resolution.
 Flow resistance was estimated by measuring the area of each outlet.
@@
 1b).
-As fully spatial relative pressure estimation methods, PPE and STE could be compared directly with $p_{\text{CFD}}$.
+PPE and STE produce spatially comprehensive relative pressure estimations ($p_{\text{STE}}$, $p_{\text{PPE}}$) that were compared directly with $p_{\text{CFD}}$.
 Relative error was evaluated at each timestep with (11) and across the entire time domain with (12).
@@
 $\varepsilon_{\text{rel}}=\frac{\left\lVert p_{\text{CFD}}-\hat{p}\right\rVert_{0,2}}{\left\lVert p_{\text{CFD}}\right\rVert_{0,2}}$ (12)
-RMSE was also calculated over the entire spatial domain at each timestep and across the entire spatiotemporal domain. In addition to these error metrics, linear regression plots were also generated.
+In addition to the relative error metrics, linear regression plots were generated.
 vWERP pressure drop estimates were compared directly with pressure drops selected from $p_{\text{CFD}}$.
@@
 2.3.3 Noise Sensitivity Analysis
-To assess the impact of estimate accuracy, noise fields [31] were added to the projected velocity fields to evaluate the performance of PPE, STE, and vWERP.
-The noise addition methodology was as follows: Velocity fields were first converted to phase data by specifying a synthetic VENC that was 1.1 times the maximum velocity component across the entire time-resolved velocity field.
+To assess the impact of noise on relative pressure estimate accuracy, random noise fields [31] were added to projected CFD velocity fields before evaluating the performance of PPE, STE, and vWERP.
+Noise was added as follows: Velocity fields were first converted to phase data by specifying a synthetic VENC that was 1.1 times the maximum velocity component across the entire time-resolved velocity field.
 Synthetic magnitude data was created by multiplying the voxelized CFD mask by a non-zero constant (leaving the non-fluid region equal to 0).
@@
 Zero-mean, complex Gaussian noise was added to k-space.
-The variance of the complex Gaussian was defined to achieve a specified SNR using the equation $\text{SNR}_{\text{db}}=10\log{\frac{P_{x}}{P_{n}}}$ ), where $P_{x}$ is the power of the signal and $P_{n}$ is the power of the noise (which is equivalent to the variance for white Gaussian noise).
-Finally, the k-space data was transformed back to image space via the inverse Fourier Transform, and the phase images were converted back to velocity images using the previously specified VENC.
+The variance of the complex Gaussian was defined to achieve a specified SNR using the equation: $\text{SNR}_{\text{db}}=10\log{\frac{P_{x}}{P_{n}}}$ ), where $P_{x}$ is the power of the signal and $P_{n}$ is the power of the noise (which is equivalent to the variance for white Gaussian noise).
+Finally, the k-space data was transformed back to image space via the inverse Fourier Transform, and the phase images were converted back to velocity images using the specified VENC.
 Twenty-five realizations of SNR = 10 noise fields and SNR = 30 noise fields (Fig.
@@
 Finally, pressure drops were evaluated for their robustness to noise with (13) and (14).
+Equations (13) and (14) are referred to in the text as “deviations” because they are designed to measure the spread of the noisy pressure estimations from the baseline (no noise) estimations.
 $e_{\Sigma}=\frac{\left\lVert\Delta p_{\text{original}}(t)-\Delta p_{\text{corrupted}}(t)\right\rVert_{0,1}}{\left\lVert\Delta p_{\text{original}}(t)\right\rVert_{0,1}}$ (13)
@@
 2.4 Flow Phantoms
-PPE, STE, and vWERP were validated with an in vitro flow setup [20]. 4D flow MRI was acquired, and twelve pressure catheter measurements per TBAD tear variation served as ground truth.
+PPE, STE, and vWERP were validated with an in vitro flow setup [20]. 4D-flow MRI was acquired, and twelve pressure catheter measurements per TBAD tear variation served as ground truth.
 2.4.1 Flow Setup
 A full description of the flow phantom can be found in [20] (pump model, exact flow specifications, etc.).
-A brief overview is provided here.
-The flow phantom geometry was designed as an embedded compliant 3D-printed TBAD derived from a clinical CT exam.
+In brief, the flow phantom geometry was designed as an embedded compliant 3D-printed TBAD model derived from a clinical CT exam.
 Physiologically accurate pulsatile flow and pressure boundary conditions were applied.
-4D flow MRI (3T; voxel size = $1.5\times 1.5\times 1.5$ mm; temporal resolution = 50 ms; VENC = 1.2–1.7 m/s) was acquired across three phantom variations: one baseline phantom (TBADOR) and two additional variations with identical anatomy but with either a reduced size FL entry tear (TBADENT) or a reduced size FL exit tear (TBADEXT) Fig.
+4D-flow MRI (3T; voxel size = $1.5\times 1.5\times 1.5$ mm; temporal resolution = 50 ms; VENC = 1.2–1.7 m/s) was acquired across three phantom variations: one baseline phantom (TBADOR) and two additional variations with identical anatomy but with either a reduced size FL entry tear (TBADENT) or a reduced size FL exit tear (TBADEXT) Fig.
 2bc.
@@
 As with the CFD model, the velocity fields were corrupted prior to pressure estimation in order to evaluate each method’s robustness to noise.
-Noise, however, is already present in the acquired 4D flow, so any additional perturbations are referred to as “corruptions”.
+Noise, however, is already present in the acquired 4D-flow, so any additional perturbations are referred to as “corruptions”.
 Each velocity component ($u_{x},u_{y},u_{z}$) was corrupted with a zero-mean, truncated Gaussian [32] with standard deviation dependent on the level of additional corruption desired ($\Delta\text{SNR}$) according to equation (15).
@@
 This included measures of mean deviation (13) and max deviation (14).
-Error averages, minimums, and maximums were recorded across each of the 50 realization combos (e.g., $\text{TBAD}_{\text{ENT}}$ at $\Delta$SNR = 10) and are reported in the supplement.
+Error averages, minimums, and maximums were recorded across each of the 50 realization combinations (e.g., $\text{TBAD}_{\text{ENT}}$ at $\Delta$SNR = 10).
 2.5 Patient Cases
@@
 3ab) to assess the clinical feasibility of pressure estimation methods and to qualitatively evaluate pressure trends against what is known in the literature, including our knowledge of FL growth rate [33].
-4D flow was acquired on 3T MRI scanners (MR750, GE Medical Systems, Milwaukee, WI, USA; Ingenia, Philips, Best, Netherlands) with the following scan parameters: voxel size = ${\sim}1.5\text{ mm}\times 1.5\text{ mm}\times 2.5\text{ mm}$, average temporal resolution = 47 ms, VENC = 2.0 m/s.
+4D-flow was acquired on 3T MRI scanners (MR750, GE Medical Systems, Milwaukee, WI, USA; Ingenia, Philips, Best, Netherlands) with the following scan parameters: voxel size = ${\sim}1.5\text{ mm}\times 1.5\text{ mm}\times 2.5\text{ mm}$, average temporal resolution = 47 ms, VENC = 2.0 m/s.
 Agreement between the pressure estimators was also evaluated.
-4D flow was acquired from a previous IRB-approved study [34] (HUM00120679).
-The data was segmented by thresholding patient MRA data and subsequently interpolating the resulting segmentation onto the 4D flow image space (Fig.
+4D-flow was acquired from a previous IRB-approved study [34] (HUM00120679).
+The data was segmented by thresholding patient MRA data and subsequently interpolating the resulting segmentation onto the 4D-flow image space (Fig.
 3cd).
-Inlet and outlet planes were generated in a 4D flow MATLAB toolbox [35].
+Inlet and outlet planes were generated in a 4D-flow MATLAB toolbox [35].
 FL GR was measured clinically as the change in maximum aortic diameter.
@@
 (b) Low-growth patient case MRA with 3D segmentation overlaid.
-(c) High-growth patient case 4D flow magnitude with velocity streamlines at peak systole overlaid.
-(d) Low-growth patient case 4D flow magnitude with velocity streamlines at peak systole overlaid.
+(c) High-growth patient case 4D-flow magnitude with velocity streamlines at peak systole overlaid.
+(d) Low-growth patient case 4D-flow magnitude with velocity streamlines at peak systole overlaid.
 3 Results
@@
 3.1.1 Spatiotemporal Analysis
-Although the spatiotemporal analysis was performed on nine unique spatiotemporal resolutions, only the 1.5 mm $\times$ 40 ms results are visualized in Fig.
+The spatiotemporal analysis was performed on nine unique spatiotemporal resolutions.
+The 1.5 mm $\times$ 40 ms results are visualized in Fig.
 4.
-Additional correlation plots and pressure drop visualizations can be found in section 1 of the supplement.
 Fig.
@@
 Fig.
-4f shows the correlation between the pressure fields calculated by STE, PPE, versus the CFD ground truth.
+4f shows the correlation between the pressure fields calculated by STE and PPE versus the CFD ground truth.
 The estimated pressure fields were highly consistent — both $p_{\text{STE}}$ and $p_{\text{PPE}}$ were very strongly correlated with $p_{\text{CFD}}$.
-PPE consistently mildly underestimated the ground truth pressure compared to STE.
+PPE consistently slightly underestimated the ground truth pressure compared to STE.
 Fig.
-4e (top) shows the pressure drop from the aortic inlet to the TL outlet as calculated by vWERP, STE, and PPE.
+4d shows the pressure drop from the aortic inlet to the TL outlet as calculated by vWERP, STE, and PPE.
 Table 2 summarizes the pressure drop linear regression results for the spatiotemporal analysis.
-All correlation coefficients were near unity for the 1.5 mm and 2.0 mm images but degraded dramatically for the 3.0 mm images.
-PPE correlation coefficients decreased the most dramatically for the 3.0 mm cases.
-Additional press drop linear regression plots can be found in subsection 1.5 of the supplement.
+All correlation coefficients were near unity for the 1.5 mm and 2.0 mm images, but degraded substantially for the 3.0 mm images.
+PPE correlation coefficients decreased the most for the 3.0 mm cases.
 Figure 4: Overview of the in silico 1.5 mm $\times$ 40 ms results.
@@
 (b) STE pressure field at peak systole.
-Aortic inlet (i) and true lumen outlet (TL4) are marked with arrow for subfigure d.
+Aortic inlet (i) and true lumen outlet (TL4) are marked with arrows for subfigure d.
 (c) PPE pressure field at peak systole (d) Pressure drop from aortic inlet to TL outlet.
-(e) Linear regression of same pressure drop.
+(e) Linear regression of the same pressure drop.
 (f) Linear regressions performed on STE and PPE pressure fields across all timesteps.
@@
 The range of total relative errors (12) is shown in Table 3.
-Averaged total relative errors can be found in subsection 1.3 of the supplement.
-Relative errors for STE were lower than for PPE; however, STE generally had larger ranges of errors, indicating that it was more sensitive to noise.
+Averaged total relative errors stayed roughly the same as noise was increased; the maximum change in average total relative error was 1.30% for PPE 2.0 mm $\times$ 60 ms when increasing from SNR=inf to SNR=10.
+Relative errors for STE were lower than for PPE for the 1.5 mm and 3.0 mm resolutions; however, STE generally had larger error ranges, indicating that it was more sensitive to noise.
 Table 3: In Silico Total Relative Error ($\varepsilon_{\text{rel}}$) Range for Sensitivity Analysis
@@
 3.0 mm 23.97% 28.18% 30.06% 15.42% 16.99% 28.98%
-In general, all methods were resilient to noise; only the 3.0 mm pressure traces in supplement subsection 1.6 had a noticeable spread of pressure values due to corrupted velocity fields.
-Furthermore, the calculated deviation values indicated that the amount of deviation from the baseline (no noise) cases increased substantially for the 3.0 mm cases.
+In general, all methods were resilient to noise; only the 3.0 mm pressure traces had a noticeable spread of pressure values due to noisy velocity fields.
+Furthermore, the calculated deviation values indicated that the spread from the baseline (no noise) cases increased substantially for the 3.0 mm cases.
 Across all resolutions and methods, the deviation values for the 1.5 mm and 2.0 mm cases remained below 4% in almost all cases and frequently below 2%.
@@
 Figure 6: Overview of in vitro flow phantom analysis.
-(a) TBADENT velocity field at peak systole.
-(b) TBADENT STE pressure field at peak systole.
-(c) TBADENT PPE pressure field at peak systole.
-(d) Linear regression across all models and planes with no corruption.
-(e) Example pressure trace over the cardiac cycle.
+(a) In vitro model STE pressure fields at peak systole.
+PPE pressure fields are not shown, but were nearly identical.
+(b) Linear regression across all models and planes with no corruption.
+(c) Example pressure trace over the cardiac cycle.
 The pressure drop was calculated from the model inlet to the model outlet (i to o in Fig.
 2).
-The “x” markers indicate where 4D flow timeframes were present.
-(f) Correlation heatmaps for each method.
-Linear regressions were run over all 50 $\Delta$SNR=10 corruption realizations per TBAD tear variation for a total of 150 unique datasets per correlation heatmap.
+The “x” markers indicate where 4D-flow timeframes were present.
+(d) Correlation heatmaps for each method.
+Linear regressions were run over all 50 $\Delta$SNR=30 corruption realizations per TBAD tear variation for a total of 150 unique datasets per correlation heatmap.
 The heatmaps indicate the distribution of pressure values in each 2D bin.
@@
 The 1D histogram above the heatmap correlation plots illustrates the number of points present within each column.
+(e) Same analysis as in d, but with $\Delta$SNR=10
 3.2.1 Validation of Pressure Estimators
 A broad overview of the flow phantom analysis results is given in Fig.
-6, with TBADENT chosen as the representative TBAD tear variation.
-More extensive results, including those for TBADOR and TBADENT, can be found in the supplement.
-Qualitatively, the pressure fields estimated by STE and PPE demonstrated the features one would expect from observing the TBADENT velocity field (Fig.
-6a).
-For instance, the high pressure impingement zone immediately following the entry tear in Fig.
-6bc closely resembles the impingement zone in the corresponding FSI model of TBADENT, shown in Fig.
-4c of Zimmermann et al.
-[20].
-Both the STE (Fig.
-6b) and PPE (Fig.
-6c) pressure fields exhibited a large pressure drop across the FL entry tear, an increase in pressure where the FL entry tear jet impinged on the wall of the FL, and higher pressure drops from the aortic inlet (i) to each TL measurement point (TL 1–5) compared to each FL measurement point (FL 1–5).
-The TBADOR and TBADEXT pressure fields were similarly coherent with our expectations.
-Fig.
-6df includes data from all TBAD tear variations.
-Across all TBAD tear variations, vWERP, STE, and PPE were strongly correlated ($r^{2}=0.91$) with catheter measurements (Fig.
-6d), with a tendency to underestimate peak pressures (Fig.
-6e).
-This trend holds across all pressure traces, which can be explored in further detail in subsection 2.3 of the supplement.
+6.
+Qualitatively, the pressure fields estimated by STE and PPE demonstrated the expected features relative to each model variation’s geometries and resulting velocity fields (Fig.
+2).
+For instance, the high pressure impingement zone immediately following the TBADENT entry tear in Fig.
+6a closely resembles the impingement zone in the corresponding FSI model of TBADENT, shown in Fig.
+4c of Zimmermann et al[20].
+Both the STE and PPE pressure fields exhibited a large pressure drop across the FL entry tear and higher pressure drops from the aortic inlet (i) to each TL measurement point (TL 1–5) compared to each FL measurement point (FL 1–5).
+The regression analysis in Fig.
+6b indicated that vWERP was the most accurate pressure estimator when no noise was added.
+A representative pressure drop is illustrated in Fig.
+6c to demonstrate what this trend in an individual pressure drop.
+Across all TBAD tear variations, vWERP, STE, and PPE were strongly correlated ($r^{2}=0.91$) with catheter measurements, with a tendency to underestimate peak pressures.
+This trend held across all pressure traces.
 3.2.2 Corruption Analysis
 Fig.
-6f illustrates the effect of the $\Delta$SNR = 10 corruption across all 50 noise realizations for each pressure estimation method.
-Corruption of the velocity fields mildly degraded the average accuracy, while the correlations were moderately weaker ($r^{2}=0.69$).
-The pressure drop mean deviation (13) and max deviation (14) results for TBADENT are summarized in Table 4.
+6d and e illustrate the effect of the $\Delta$SNR = 30 and $\Delta$SNR = 10 corruptions across all 50 noise realizations for each pressure estimation method.
+Corruption of the velocity fields mildly degraded the average accuracy, while the correlations were moderately weaker ($r^{2}$ = 0.69–0.70).
+The pressure drop mean deviation (13) and max deviation (14) results for all models are summarized in Table 4.
 For the FL and TL rows, deviations were averaged across all numbered pressure drops for that lumen (e.g., aortic inlet to TL1, aortic inlet to TL2, etc.).
-Additional tables for TBADOR and TBADEXT can be found in the supplement in subsection 2.2, and a more in-depth analysis for each pressure trace and the range of estimations can be found in the supplement subsection 2.3.
 As expected, the average deviation increased with noise.
 Average deviations were also higher in the true lumen across all models.
-PPE was significantly more resilient to corruption compared to vWERP and STE — the average PPE deviations were roughly half of the other method’s deviations.
-Table 4: In Vitro Average Deviations Across TBADENT
+PPE was significantly more resilient to corruption compared to vWERP and STE — the average PPE deviations were roughly half of the other methods’ deviations.
+Table 4: In Vitro Average Deviations Across All Models
 vWERP STE PPE
-SNR = 30 SNR = 10 SNR = 30 SNR = 10 SNR = 30 SNR = 10
+TBADOR SNR = 30 SNR = 10 SNR = 30 SNR = 10 SNR = 30 SNR = 10
+$e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$
+FL 1.30% 1.06% 5.00% 3.60% 0.97% 0.71% 3.58% 2.44% 0.48% 0.36% 1.54% 1.16%
+TL 2.13% 1.46% 8.31% 5.22% 1.83% 1.26% 6.97% 4.44% 0.95% 0.69% 3.08% 2.14%
+Outlet 2.09% 1.51% 7.27% 5.25% 0.87% 0.62% 3.29% 2.22% 0.42% 0.30% 1.38% 0.97%
+TBADENT SNR = 30 SNR = 10 SNR = 30 SNR = 10 SNR = 30 SNR = 10
 $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$
@@
 Outlet 2.38% 2.50% 8.55% 8.25% 1.22% 1.16% 4.73% 4.15% 0.58% 0.57% 1.88% 1.75%
+TBADEXT SNR = 30 SNR = 10 SNR = 30 SNR = 10 SNR = 30 SNR = 10
+$e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$ $e_{\Sigma}$ $e_{\max}$
+FL 0.72% 0.57% 2.58% 1.95% 1.37% 1.08% 5.59% 3.88% 0.69% 0.55% 2.27% 1.77%
+TL 1.49% 1.16% 5.03% 3.71% 2.73% 2.10% 10.58% 7.40% 1.35% 1.12% 4.42% 3.55%
+Outlet 0.61% 0.60% 2.02% 1.94% 1.18% 1.12% 4.43% 3.88% 0.54% 0.54% 1.79% 1.69%
 3.3 Patient Cases
-For all eight patient cases, PPE and STE produced pressure fields that were coherent with our expectations — relative pressure was highest in the aortic arch and steadily decreased when moving distally.
-Pressure drops were visibly higher in the TL, where most of the net flow was present [36].
+For all eight patient cases, PPE and STE produced pressure fields that were consistent with our expectations — relative pressure was highest in the aortic root and decreased when moving distally.
+Pressure drops were visibly higher in the TL, where a majority of the net flow was present [36].
 Pressure drops in the FL were generally lower than in the TL, although the magnitude of this trend was dependent on the calculated growth rate of each patient [33].
 As expected, lower FL growth rate patients had a higher pressure drop across the FL, and higher FL growth rate patients had a lower pressure drop across the FL (Fig.
-7ab).
+7abcd).
 The PPE, STE, and vWERP pressure drops were analyzed for agreement in Fig.
-7cd, and additional comparisons can be found in section 3 of the supplement.
+7efg.
 Correlations between all pressure estimation methods were extremely strong ($r^{2}$ = 0.97–0.98).
@@
 This ordering of the results matched the outcome of the in vitro phantom analysis, where vWERP also produced the highest pressure drops.
-Standard practice was to run the patient cases with nearest-neighbor upsampling by a factor of 2.
+Standard practice was to run the patient cases with nearest-neighbor upsampling by a factor of two.
 At this level, pressure estimations were completed in a matter of minutes as timed by MATLAB’s timeit function.
@@
 (b) STE pressure field of a low-growth patient case at peak systole.
-(c) TL and FL correlation plot of STE vs.
+(c) STE pressure field of a second high-growth patient case at peak systole.
+(d) STE pressure field of a second low-growth patient case at peak systole.
+(e) TL and FL correlation plot of STE vs.
 vWERP across all patient cases.
-(d) TL and FL correlation plot of PPE vs.
+(f) TL and FL correlation plot of PPE vs.
 vWERP across all patient cases.
+(g) Fully spatial correlation plot of PPE vs.
+STE across all patient cases.
 4 Discussion
@@
 The in vitro analysis was performed on a compliant TBAD flow phantom with three unique entry and exit tear variants.
-Twelve pressure catheter measurements were made in each flow phantom tear variation, combining the benefits of acquired 4D flow and significantly more ground truth pressure measurements than would be allowed clinically.
+Twelve pressure catheter measurements were made in each TBAD flow phantom tear variation, combining the benefits of acquired 4D-flow and significantly more ground truth pressure measurements than would be allowed clinically.
 The in vivo analysis was performed on eight patient cases.
@@
 Because the rest of the model proximal to the narrow TL was disrupted by the same constant, the four FL pressure drops we probed remained accurate.
-In the patient cases, small segments of 2–8 voxels with erroneous velocity measurements led to small regions (2–8 voxels as well) with massive pressure errors.
+In the patient cases, small segments of 2–8 voxels with erroneous velocity measurements led to small regions (2–8 voxels as well) with large pressure errors.
 High (but localized) erroneous velocity gradients led to corresponding localized pressure errors.
-In response to these findings, care was taken to pre-process patient data to remove 4D flow voxels with magnitudes/orientations significantly different from their neighbors.
+In response to these findings, care was taken to pre-process patient data to remove 4D-flow voxels with magnitudes/orientations significantly different from their neighbors.
 If STE and PPE are used clinically, care must be taken to appropriately filter out erroneous velocity measurements to ensure errors are not magnified in the corresponding relative pressure fields.
-Errors generally increased when moving proximally from the TL outlet, which is expected given that the relative pressure fields were shifted in reference to the TL outlet pressure.
+Errors generally increased when moving proximally from the TL outlet, which is expected, given that the relative pressure fields were shifted in reference to the TL outlet pressure.
 Figure 8: Overview of in silico origin shift analysis for STE.
-(a) Overview of origin shift method for one voxel in 2D.
+(a) Overview of the origin shift method for one voxel in 2D.
 Each square indicates a different shift in the origin by a subvoxel amount, forcing a different interpolation.
@@
 Increased spatial resolution generally increased the accuracy of the estimated pressure fields.
-Accuracy dipped when temporal resolution decreased mainly because the pressure estimators did not have the time frames available to capture the highly transient pressure drop at peak systole.
+Accuracy decreased when temporal resolution decreased, mainly because the pressure estimators did not have the time frames available to capture the highly transient pressure drop at peak systole.
 Errors were primarily caused by underestimating pressure extrema (both systolic peaks and diastolic valleys) and a temporal shift in the latter half of the cardiac cycle.
-This behavior was more drastic in the 4D flow MRI acquisitions compared to the synthetic 4D flow images.
-The synthetic 4D flow images were constructed with instantaneous velocity measurements, which meant there was no temporal averaging.
-Temporal averaging, of course, occurred in the 4D flow scans.
+This behavior was more drastic in the 4D-flow MRI acquisitions compared to the synthetic 4D-flow images.
+The synthetic 4D-flow images were constructed with instantaneous velocity measurements, which meant there was no temporal averaging.
+Temporal averaging, of course, occurred in the 4D-flow scans.
 In the in silico analysis, the estimators may have missed peak pressures, but they accurately estimated the pressures based on the time frames they had access to.
-Meanwhile, the 4D flow scans both undersampled the peak systolic pressure drop and temporally averaged the time frames that were sampled, leading to more pronounced underestimation.
+Meanwhile, the 4D-flow scans both undersampled the peak systolic pressure drop and temporally averaged the time frames that were sampled, leading to more pronounced underestimation.
 4.3 Effect of Noise
@@
 Whereas higher resolution was almost always better in the SNR=inf cases, this trend did not precisely hold in the noisy cases.
-Low spatial resolution combined with high temporal resolution resulted in a higher spread in estimated pressures, although the average pressures across noise realizations were still close to the SNR=inf cases.
-Deviations from the baseline pressure drop estimation were also routinely higher in the TL across both in silico and in vitro analyses.
+Low spatial resolution combined with high temporal resolution resulted in larger deviations in estimated pressures, although the average pressures across noise realizations were still close to the SNR=inf cases.
+Deviation from the baseline relative pressure drop were also routinely higher in the TL compared to the FL across both in silico and in vitro analyses.
 4.4 Pros and Cons of Pressure Estimators
@@
 More clear differences in accuracy were present, however, in the sensitivity and spatiotemporal analyses.
-Based on our analysis, vWERP should be the preferred estimator when calculating one-dimensional pressure drops.
-It had the highest accuracy in the 4D flow MRI acquisitions and was significantly faster than STE.
+Based on our analysis, vWERP outperformed STE and PPE when calculating one-dimensional pressure drops.
+It had the highest accuracy in the 4D-flow MRI acquisitions and was significantly faster than STE.
 Although vWERP performed worse than PPE and STE in the in silico cases, vWERP’s improved performance in the in vitro cases suggests that there are aspects of MRI simulation that we are missing in our synthetic MRI generation methodology.
-When a fully 3D pressure field is desired, the ideal estimator depends on the SNR and resolution of the 4D flow acquisition.
-STE should be the standard choice (especially at lower resolutions).
-However, if a low SNR is expected, PPE may be better than STE.
+When a fully 3D pressure field is desired, the ideal estimator depends on the SNR and resolution of the 4D-flow acquisition.
+Our data suggests that STE performs better, especially at lower spatial resolutions.
+However, if a low SNR is expected, PPE may outperform STE.
 Similar to [15], STE outperformed PPE in terms of accuracy; however, PPE appeared to have a slight edge in its resistance to noise.
@@
 PPE and STE agreed strongly, and vWERP tended to estimate the largest pressure drop, mimicking the trend of the in vitro analysis.
-All pressure estimators were remarkably resilient to noise and successfully estimated the ground truth pressure, albeit with tendencies to underestimate.
+All pressure estimators were remarkably resilient to noise and successfully estimated the ground truth pressure, albeit with a tendency to underestimate peak relative pressures.
 The time to compute the pressure estimations was on the order of minutes, demonstrating the possibility of being integrated into a clinical workflow.